# EW corrections and Heavy Boson Radiation at a high-energy muon collider

### v2

@@ 1 Introduction
In recent years a novel interest for a muon collider has arisen, motivated by its great potential for the investigation of the fundamental interactions of Nature MuonCollider:2022xlm ; Aime:2022flm ; Black:2022cth ; Maltoni:2022bqs ; Belloni:2022due ; Accettura:2023ked . A key aspect of a muon collider is the possibility of accelerating elementary particles at energies of several TeV’s Delahaye:2019omf ; Bartosik:2020xwr ; Schulte:2021hgo ; Long:2020wfp ; MuonCollider:2022nsa ; MuonCollider:2022ded ; MuonCollider:2022glg , leading to the possibility to probe fundamental interactions at unprecedented energies. It offers the potential to significantly advance our understanding of fundamental particle physics, enabling in-depth studies of the Standard Model (SM) Higgs boson Han:2020pif ; Buttazzo:2020uzc ; Han:2021lnp ; Reuter:2022zuv ; Celada:2023oji ; Forslund:2022xjq ; Chen:2022yiu ; Ruhdorfer:2023uea ; Han:2023njx ; Dermisek:2023rvv ; Liu:2023yrb ; Li:2024joa ; Cassidy:2023lwd , searches for beyond-the-SM (BSM) heavy Higgs bosons Han:2022edd ; Bandyopadhyay:2020otm ; Han:2021udl ; Bandyopadhyay:2024plc ; Ouazghour:2023plc ; Jueid:2023qcf , investigations of dark matter Han:2020uak ; Capdevilla:2021fmj ; Han:2022ubw ; Cesarotti:2024rbh ; Asadi:2023csb ; Belfkir:2023vpo ; Jueid:2023zxx , constraints on lepton-universality violation Huang:2021nkl ; Huang:2021biu ; He:2024dwh , exploration of the muon $g-2$ anomaly Buttazzo:2020ibd ; Capdevilla:2020qel ; Yin:2020afe ; Capdevilla:2021rwo , and tests of a wide range of new physics scenarios Gu:2020ldn ; Costantini:2020stv ; Bandyopadhyay:2021pld ; Qian:2021ihf ; Bao:2022onq ; Ghosh:2023xbj ; Dermisek:2023tgq ; Altmannshofer:2023uci ; Liu:2023jta ; Sun:2023cuf ; Kwok:2023dck ; Cesarotti:2023sje ; Ema:2023buz ; Li:2023tbx ; Bhattacharya:2023beo ; Das:2024ekt .
@@
Finally, we scrutinise the contribution of Heavy-Boson-Radiation (HBR), in other words the emission of weak bosons $W,Z$, and $H$. We consider different scenarios, where HBR is recombined or not with particles in the final state $F$ in $\mu^{+}\mu^{-}\to F$ production. We also consider as physical objects in the final state $F$ the so-called “EW jets”, obtained by the clustering of $W$ and $Z$ bosons. We compare the impact of HBR, which is calculated exactly and taking into account phase-space cuts, with the one of NLO EW corrections, also calculated exactly, and discuss their relative size and possible cancellations. We show how HBR leads in general to much smaller contributions than their virtual counterparts, only marginally compensating for the large effects due to the latter.
@@ 2.1 The automation of NLO EW corrections
[…]
@@ 2.2.1 Amplitude level
$L(|r_{kl}|,M^{2})\equiv\frac{\alpha}{4\pi}\log^{2}{\frac{|r_{kl}|}{M^{2}}}\qquad{\rm and}\qquad l(|r_{kl}|,M^{2})\equiv\frac{\alpha}{4\pi}\log{\frac{|r_{kl}|}{M^{2}}}\,.$ ) . (2.3)
@@
Via the DP algorithm it is possible to calculate in LA one-loop EW corrections of a generic SM scattering amplitude ${\cal{M}}$, which are typically denoted as $\delta{\cal{M}}$, to the Born approximation, which is instead typically denoted as ${\cal{M}}_{0}$. For any individual helicity configuration of the amplitude ${\cal{M}}$, the DP algorithm allows to write $\delta{\cal{M}}$ as a function of the logarithms in Eq. (2.3), the couplings of each external field to the gauge bosons (and another possible field) or associated quantities such as electroweak Casimir operators, and tree-level amplitudes as ${\cal{M}}_{0}$ or similar ones with one or two of the external fields replaced by, e.g., $SU(2)$ partners w.r.t. the case of ${\cal{M}}_{0}$. This is precisely at the origin of the resurgence of the interest in EWSL and the DP algorithm in the past few years: EWSL can be computed in a much faster and more stable way than the exact NLO EW corrections and this approach can be (supposedly) extended to the BSM case, capturing the leading corrections at high energies.
@@
  • For the helicity configuration considered, in the high-energy limit, the tree-level amplitude ${\cal{M}}_{0}$ must not be mass-suppressed by powers of the form $(M_{W}/\sqrt{s})^{k}$ with $k>0$. In other words, by dimensional analysis, a $2\rightarrow n$ process requires that $[{\cal{M}}]=(\,\textrm{GeV})^{2-n}$ and therefore
@@ 3.1 Calculation set-up
The two classes of processes entail completely different kinematics, especially in the bulk of the associated cross sections. Direct production is dominated by the phase-space region $s\simeq S$, where $\sqrt{s}$ is the total energy of the partonic process in its rest frame while $\sqrt{S}$ is the collider energy. Configurations with $s\neq S$ are induced by the emissions of photons, in particular the initial-state-radiation (ISR), which is taken into account in the collinear limit directly via PDF evolution of the (anti)muon in the (anti)muon $\Gamma_{\mu^{\pm}/\mu^{\pm}}$ or otherwise via the NLO EW corrections. VBF production is instead dominated by the phase-space region $\sqrt{s}\simeq m(F)\simeq\sum_{i\in F}m_{i}$, where $i$ is any particle that is part of the final state $F$ and $m(F)$ is the invariant mass of the final state $F$. In other words, The $\mu\mu^{-}/\nu_{\mu}\bar{\nu}_{\mu}$ additional pair in the final state carries away most of the energy of the colliding muons and it typically does it along the beam pipe axis. The hard process is in fact $VV\rightarrow F$, and the leading contributions, especially at very high energies, can be simulated in the so-called Effective-Vector-boson-Approximation (EVA) Dawson:1984gx ; Kane:1984bb , which has already been implemented in MadGraph5_aMC@NLO Ruiz:2021tdt . Instead of simulating $22\rightarrow n+$ 2 processes, where $n$ is the multiplicity of $F$, $2\rightarrow n$ matrix elements are sufficient ($VV\rightarrow F$) and the “$V$ in the muon” can be modeled similarly to what is done in the case of the photon in the lepton in the Weizsäcker–Williams approximation vonWeizsacker:1934nji ; Williams:1934ad . Not only, as already mentioned Sec. 2.1, these effects can also be resummed and taken into account in a PDF formalism, as shown in Refs. Han:2020uid ; Han:2021kes ; Garosi:2023bvq . [^9] It is interesting to note that since, at least for low multiplicities, $m(F)\simeq\sum_{i\in F}m_{i}\simeq M_{W}$, not only logarithms of the form $\log(s/M_{W}^{2})$ entering the PDFs are important but also power corrections of the form $(m(F)/M_{W})^{n}$ to the matrix elements cannot be neglected. However, this aspect is beyond the scope of this paper and it will not be investigated here.
@@
$\eta(X)<2.44\,,\qquad p_{T}(X)>100\,{\,\textrm{GeV}}\,,\qquad\Delta R(X,Y)>0.4\,,\qquad\,$ (3.3)
@@
When energies of several TeV’s are reached, one may wonder if not only photons but also $H$, $Z$ and $W$ bosons have to be clustered together with heavy particles. Indeed, at such energies EW radiation is expected to lead to large effects. Their origin is the real-emission counterpart of the virtual EWSL: soft and/or collinear enhancements in the radiation. On the one hand, from an experimental point of view, if massive particles are very close in angle, their decay products are expected to be clustered in a single fat jet. On the other hand, from a theoretical point of view, real $H,Z$, and $W$ radiation induces ${\cal O}(\alpha)$ corrections to the inclusive direct production of a given final state $F$, which therefore are of the same perturbative order of NLO EW corrections or EWSL. In fact, they can be formally considered as part of the NLO EW corrections to the hard process. This aspect has already been discussed in the literature in the context of present and future hadron-colliders, e.g., in Refs. Frixione:2014qaa ; Frixione:2015zaa ; Czakon:2017wor , and this new contribution has been denoted in these references as Heavy-Boson-Radiation (HBR).
@@
$M_{H}=125\text{ GeV},\quad m_{t}=173.3\text{ GeV}.\$ (3.6)
@@ 3.2.2 EWSL and their approximate resummation
The r.h.s. of Eq. (3.20) says that if $\sigma_{\rm EXP_{EW}}$ is expanded in powers of $\alpha$ the NLO EW prediction is captured exactly, while beyond ${\cal O}(\alpha)$ the resummed tower of EWSL of order $\alpha^{n}\log^{k}(s/M_{W}^{2})$ with $n>1$ and $k=2n,2n-1$ is approximated via simple exponentiation. We stress again that we do not claim we are doing NLL resummation of EWSL. We instead want to study when and if this procedure is necessary, by comparing $\delta_{\rm NLO_{EW}}$ with the relative corrections induced by $\sigma_{\rm EXP_{EW}}$, namely
@@ 3.3 List of aspects investigated in this work
    When NLO EW corrections reach, or even worse surpass, the relative size of $-100\%$ of the LO, resummation of EWSL does not concern precision but the physical sense of predictions. We want to investigate when we should expect such situations at 3 and 10 TeV collisions.
@@ 4.1.2 Importance of logarithms involving ratios of invariants
In order to minimise the overlap with the effects discussed in the previous section, ${\rm SDK_{weak}}$ vs. ${\rm SDK_{0}}$, we consider final states with only neutrally charged particles, in particular: the $ZZ$, $ZH$, $ZZZ$, and $ZZH$ production processes. The layout of the plots in this section is very similar to those shown in the previous section. The only difference w.r.t. them is that we show here ${\rm SDK_{weak}}$ as defined in Eq. (3.14) (again displayed as a solid red line) and the same quantity where we set $\delta_{s\rightarrow r_{kl}}=0$ (solid green line) in the aforementioned equation.
[…]
@@ 4.3 Heavy Boson Radiation (HBR)
One of the widespread assumptions about a high energy muon collider is that soft and/or collinear splittings involving heavy weak bosons ($W$, $Z$ and $H$) will lead to ${\cal O}(1)$ corrections, similarly to the case of QCD at the LHC, see e.g. Han:2020uid ; Han:2021lnp . On the one hand, this implies that a muon collider can be treated as a EW boson collider, Costantini:2020stv ; Buttazzo:2020uzc ; Han:2020uid ; Han:2021lnp ; Ruiz:2021tdt ; AlAli:2021let ; Accettura:2023ked leading to very large cross sections for VBF processes. On the other hand, the emission of $W,Z$ and possibly $H$ bosons, i.e., what has been dubbed in this work as HBR, is in general expected to lead to large effects or even becoming dominant w.r.t. the case with no radiation. This dynamics has already been studied in detail in particular scenarios, e.g. showing how it can help to gain sensitivity to new short-distance physical laws Chen:2022msz .
@@ 4.3.1 Final states: $t\bar{t}$, $W^{+}W^{-}$
[…]
While the EWSL from virtual corrections receive contributions from $W$ and $Z$ bosons, at ${\cal O}(\alpha)$ the HBR receive contributions only from the $Z$ boson (and the $H$ boson) emissions and especially are suppressed due to the phase-space cuts. One cannot simply estimate the HBR rescaling the LO predictions by the EWSL as done in the case of virtual contributions. Moreover, as pointed out in Ref. Bagdatova:2024aem , the fact that the $W,Z$, and $H$ have a non-zero mass cannot be neglected and has an impact also in the results that are obtained.

### v3

@@ 1 Introduction
In recent years a novel interest for a muon collider has arisen, motivated by its great potential for the investigation of the fundamental interactions of Nature MuonCollider:2022xlm ; Aime:2022flm ; Black:2022cth ; Maltoni:2022bqs ; Belloni:2022due ; Accettura:2023ked . A key aspect of a muon collider is the possibility of accelerating elementary particles at energies of several TeV’s Delahaye:2019omf ; Bartosik:2020xwr ; Schulte:2021hgo ; Long:2020wfp ; MuonCollider:2022nsa ; MuonCollider:2022ded ; MuonCollider:2022glg , leading to the possibility to probe fundamental interactions at unprecedented energies. It offers the potential to significantly advance our understanding of fundamental particle physics, enabling in-depth studies of the Standard Model (SM) Higgs boson Han:2020pif ; Buttazzo:2020uzc ; Han:2021lnp ; Reuter:2022zuv ; Celada:2023oji ; Forslund:2022xjq ; Chen:2022yiu ; Ruhdorfer:2023uea ; Han:2023njx ; Dermisek:2023rvv ; Liu:2023yrb ; Li:2024joa ; Cassidy:2023lwd , searches for beyond-the-SM (BSM) heavy Higgs bosons Han:2022edd ; Bandyopadhyay:2020otm ; Han:2021udl ; Bandyopadhyay:2024plc ; Ouazghour:2023plc ; Jueid:2023qcf , investigations of dark matter Han:2020uak ; Capdevilla:2021fmj ; Han:2022ubw ; Cesarotti:2024rbh ; Asadi:2023csb ; Belfkir:2023vpo ; Jueid:2023zxx , constraints on lepton-universality violation Huang:2021nkl ; Huang:2021biu ; He:2024dwh , exploration of the muon g2g-2 anomaly Buttazzo:2020ibd ; Capdevilla:2020qel ; Yin:2020afe ; Capdevilla:2021rwo , and tests of a wide range of new physics scenarios Gu:2020ldn ; Costantini:2020stv ; Bandyopadhyay:2021pld ; Qian:2021ihf ; Bao:2022onq ; Ghosh:2023xbj ; Dermisek:2023tgq ; Altmannshofer:2023uci ; Liu:2023jta ; Sun:2023cuf ; Kwok:2023dck ; Cesarotti:2023sje ; Ema:2023buz ; Li:2023tbx ; Bhattacharya:2023beo ; Das:2024ekt .
@@
Finally, we scrutinise the contribution of Heavy-Boson-Radiation (HBR), in other words the emission of weak bosons $W,Z$, and $H$. We consider different scenarios, where HBR is recombined or not with particles in the final state $F$ in $\mu^{+}\mu^{-}\to F$ production. We also consider as physical objects in the final state $F$ the so-called “EW jets”, obtained by the clustering of $W$ and ZZ bosons. We compare the impact of HBR, which is calculated exactly and taking into account phase-space cuts, with the one of NLO EW corrections, also calculated exactly, and discuss their relative size and possible cancellations. We show how HBR leads in general to much smaller contributions than their virtual counterparts, only marginally compensating for the large effects due to the latter.
@@ 2.1 The automation of NLO EW corrections
one can easily see that the bulk of the cross section comes from regions where $x\to 1$. [^5] We clearly assume that there are no other enhancement effects, such as the direct production of a new resonant heavy state. The peculiar dependence of the leptonic PDFs required some changes in MadGraph5_aMC@NLO, in order to have an efficient numerical integration: the first, trivial, is to flatten out the integrable divergence via a suitable change of integration variables Frixione:2021zdp . The second one specific to the computation of NLO EW corrections, is to devise a phase-space mapping where the event and its soft/collinear counterterms are evaluated at the same values of the Bjorken $x$. Such a new mapping is documented in the appendix of Ref. Bertone:2022ktl . Finally, the existence of NLL densities in different factorisation schemes requires the inclusion in the short-distance cross section of additional terms (finite contributions to the initial-state counterterms). The same applies to the case when LL-accurate PDFs are employed, in order to attain formal NLO accuracy.
@@ 2.2.1 Amplitude level
[…]
@@
[…]
@@
  • For the helicity configuration considered, in the high-energy limit, the tree-level amplitude ${\cal{M}}_{0}$ must not be mass-suppressed by powers of the form $(M_{W}/\sqrt{s})^{k}$ with $k>0$. In other words, by dimensional analysis, a $2\rightarrow n$ process requires that ${\cal{M}}]=(\,\textrm{GeV})^{2-n}$ and therefore
@@ 3.1 Calculation set-up
The two classes of processes entail completely different kinematics, especially in the bulk of the associated cross sections. Direct production is dominated by the phase-space region $s\simeq S$, where $\sqrt{s}$ is the total energy of the partonic process in its rest frame while $\sqrt{S}$ is the collider energy. Configurations with $s\neq S$ are induced by the emissions of photons, in particular the initial-state-radiation (ISR), which is taken into account in the collinear limit directly via PDF evolution of the (anti)muon in the (anti)muon $\Gamma_{\mu^{\pm}/\mu^{\pm}}$ or otherwise via the NLO EW corrections. VBF production is instead dominated by the phase-space region $\sqrt{s}\simeq m(F)\simeq\sum_{i\in F}m_{i}$, where $i$ is any particle that is part of the final state $F$ and $m(F)$ is the invariant mass of the final state $F$. In other words, The $\mu\mu^{-}/\nu_{\mu}\bar{\nu}_{\mu}$ additional pair in the final state carries away most of the energy of the colliding muons and it typically does it along the beam pipe axis. The hard process is in fact $VV\rightarrow F$, and the leading contributions, especially at very high energies, can be simulated in the so-called Effective-Vector-boson-Approximation (EVA) Dawson:1984gx ; Kane:1984bb , which has already been implemented in MadGraph5_aMC@NLO Ruiz:2021tdt . Instead of simulating $2\rightarrow n+2$ processes, where $n$ is the multiplicity of $F$, $2\rightarrow n$ matrix elements are sufficient ($VV\rightarrow F$) and the “$V$ in the muon” can be modeled similarly to what is done in the case of the photon in the lepton in the Weizsäcker–Williams approximation vonWeizsacker:1934nji ; Williams:1934ad . Not only, as already mentioned Sec. 2.1, these effects can also be resummed and taken into account in a PDF formalism, as shown in Refs. Han:2020uid ; Han:2021kes ; Garosi:2023bvq . [^9] It is interesting to note that since, at least for low multiplicities, $m(F)\simeq\sum_{i\in F}m_{i}\simeq M_{W}$, not only logarithms of the form $\log(s/M_{W}^{2})$ entering the PDFs are important but also power corrections of the form $(m(F)/M_{W})^{n}$ to the matrix elements cannot be neglected. However, this aspect is beyond the scope of this paper and it will not be investigated here.
@@
$\eta(X)<2.44\,,\qquad p_{T}(X)>100\,{\,\textrm{GeV}}\,,\qquad\Delta R(X,Y)>0.4\,$$\qquad\,$ (3.3)
@@
When energies of several TeV’s are reached, one may wonder if not only photons but also $H$, ZZ and $W$ bosons have to be clustered together with heavy particles. Indeed, at such energies EW radiation is expected to lead to large effects. Their origin is the real-emission counterpart of the virtual EWSL: soft and/or collinear enhancements in the radiation. On the one hand, from an experimental point of view, if massive particles are very close in angle, their decay products are expected to be clustered in a single fat jet. On the other hand, from a theoretical point of view, real $H,Z$, and $W$ radiation induces ${\cal O}(\alpha)$ corrections to the inclusive direct production of a given final state $F$, which therefore are of the same perturbative order of NLO EW corrections or EWSL. In fact, they can be formally considered as part of the NLO EW corrections to the hard process. This aspect has already been discussed in the literature in the context of present and future hadron-colliders, e.g., in Refs. Frixione:2014qaa ; Frixione:2015zaa ; Czakon:2017wor , and this new contribution has been denoted in these references as Heavy-Boson-Radiation (HBR).
@@
$M_{H}=125\text{ GeV},\quad m_{t}=173.3\text{ GeV}.\\$ (3.6)
@@ 3.2.2 EWSL and their approximate resummation
The r.h.s. of Eq. (3.20) says that if $\sigma_{\rm EXP_{EW}}$ is expanded in powers of $\alpha$ the NLO EW prediction is captured exactly, while beyond ${\cal O}(\alpha)$ the resummed tower of EWSL of order $\alpha^{n}\log^{k}(s/M_{W}^{2})$ with $n>1$ and k=2n,2n1k=2n,2n-1 is approximated via simple exponentiation. We stress again that we do not claim we are doing NLL resummation of EWSL. We instead want to study when and if this procedure is necessary, by comparing $\delta_{\rm NLO_{EW}}$ with the relative corrections induced by $\sigma_{\rm EXP_{EW}}$, namely
@@ 3.3 List of aspects investigated in this work
    When NLO EW corrections reach, or even worse surpass, the relative size of $100\%$ of the LO, resummation of EWSL does not concern precision but the physical sense of predictions. We want to investigate when we should expect such situations at 3 and 10 TeV collisions.
@@ 4.1.2 Importance of logarithms involving ratios of invariants
In order to minimise the overlap with the effects discussed in the previous section, $\rm SDK_{weak}$} vs. $\rm SDK_{0}$}, we consider final states with only neutrally charged particles, in particular: the $ZZ$, ZHZH, ZZZZZZ, and $ZZH$ production processes. The layout of the plots in this section is very similar to those shown in the previous section. The only difference w.r.t. them is that we show here $\rm SDK_{weak}$} as defined in Eq. (3.14) (again displayed as a solid red line) and the same quantity where we set $\delta_{s\rightarrow r_{kl}}=0$ (solid green line) in the aforementioned equation.
We start by showing again the same observables considered for $ZZ$ and $ZH$ production in the previous section, $p_{T}(Z_{2})$ for $ZZ$ production in Fig. 6 and $p_{T}(H)$ for $ZH$ production in Fig. 7. In Fig. 6 we clearly see that for $p_{T}(Z_{2})\gtrsim 250~\,\textrm{GeV}$ ($p_{T}(Z_{2})\gtrsim 800~\,\textrm{GeV}$) at 3 TeV (10 TeV) collisions, the very good accuracy of the $\rm SDK_{weak}$} prediction ($\delta_{{\rm SDK_{weak}}}-\delta_{\rm NLO_{EW}}$ constant and of ${\cal O}(1\%)$) is much degraded when $\delta_{s\rightarrow r_{kl}}=0$, i.e. the green line in the second inset. Indeed, while at large $p_{T}(Z_{2})$ we see $\delta_{{\rm SDK_{weak}}}-\delta_{\rm NLO_{EW}}\simeq 1\%$, at smaller values, $p_{T}(Z_{2})\simeq 250~\,\textrm{GeV}$ ($p_{T}(Z_{2})\simeq 800~\,\textrm{GeV}$) at 3 TeV (10 TeV), we notice that $\delta_{{\rm SDK_{weak}}}-\delta_{\rm NLO_{EW}}\simeq 30\%$. Since the $p_{T}(Z_{2})$ distribution is quite flat, this discrepancy at small $p_{T}(Z_{2})$ has an effect also at the level of the total cross section; setting $\delta_{s\rightarrow r_{kl}}=0$ we find that, for both 3 and 10 TeV, with the cuts considered $\delta_{{\rm SDK_{weak}}}-\delta_{\rm NLO_{EW}}\simeq 10\%$ also for the total cross section. A similar (quite constant) discrepancy is observed in the $m(ZZ)$ distribution, too. These results are clearly dependent on the cuts in (3.3), in particular the one on the pseudorapidity of the $Z$ bosons. Setting $\delta_{s\rightarrow r_{kl}}=0$ logarithms of the form, e.g., $L(|t|,s)$ are omitted and, in the proximity of the pseudorapidity cuts, such logarithms are of the same order for both the 3 and 10 TeV results.
@@ 4.3 Heavy Boson Radiation (HBR)
One of the widespread assumptions about a high energy muon collider is that soft and/or collinear splittings involving heavy weak bosons ($W$, ZZ and $H$) will lead to ${\cal O}(1)$ corrections, similarly to the case of QCD at the LHC, see e.g. Han:2020uid ; Han:2021lnp . On the one hand, this implies that a muon collider can be treated as a EW boson collider, Costantini:2020stv ; Buttazzo:2020uzc ; Han:2020uid ; Han:2021lnp ; Ruiz:2021tdt ; AlAli:2021let ; Accettura:2023ked leading to very large cross sections for VBF processes. On the other hand, the emission of $W,Z$ and possibly $H$ bosons, i.e., what has been dubbed in this work as HBR, is in general expected to lead to large effects or even becoming dominant w.r.t. the case with no radiation. This dynamics has already been studied in detail in particular scenarios, e.g. showing how it can help to gain sensitivity to new short-distance physical laws Chen:2022msz .
@@ 4.3.1 Final states: $t\bar{t}$, $W^{+}W^{-}$
Let us see how it is possible that the contribution from HBR is much smaller than the one from $\rm NLO_{EW}$, although $\sqrt{S}\gg M_{W}$. First of all, in order to select the direct-production mechanism and exclude VBF configuration, the cut in (3.1) is present in our simulations. Moreover, also the cuts in (3.3) are present and have to be satisfied. Thus, the phase space of the $H$ and ZZ radiation is much more constrained than one could naively expect. As a consequence, the double logarithms of the form $L(S,M^{2}_{B})$, as those appearing in the virtual contributions, cannot be present. A larger contribution can be observed if the cut in Eq. (3.1) is relaxed to $m(t\bar{t})\geq 0.5\,\sqrt{S}$, as shown in Fig. 19. However, the effects from HBR are far from being of ${\cal O}(1)$ and are especially much smaller than NLO EW corrections.
While the EWSL from virtual corrections receive contributions from $W$ and ZZ bosons, at ${\cal O}(\alpha)$ the HBR receive contributions only from the $Z$ boson (and the $H$ boson) emissions and especially are suppressed due to the phase-space cuts. One cannot simply estimate the HBR rescaling the LO predictions by the EWSL as done in the case of virtual contributions. Moreover, as pointed out in Ref. Bagdatova:2024aem , the fact that the $W,Z$, and $H$ have a non-zero mass cannot be neglected and has an impact also in the results that are obtained.